OpenStudy (anonymous):

\[sec\left(cos^{-1}\frac 12\right)\] \[cos\alpha=\frac 12\] \[\alpha=\frac {\pi} 3\] where from here?

OpenStudy (anonymous):

oh is it just \[sec(\frac \pi 3)\]

zepdrix (zepdrix):

\[\large \sec\left(\arccos\frac{1}{2}\right) \qquad = \qquad \sec\left(\frac{\pi}{3}\right) \qquad = \qquad ?\] Yah I think you've got the right idea :)

OpenStudy (anonymous):

2

zepdrix (zepdrix):

Mmmmm yah sounds good c:

OpenStudy (anonymous):

yay....let me try another one

zepdrix (zepdrix):

Make sure you're comfortable with this notation \(\large \cos^{-1}x\). The -1 is never a power. It always represents the inverse function when we're dealing with trig.

zepdrix (zepdrix):

\[\large \cos^{-1}x \neq \frac{1}{\cos x}\]

OpenStudy (anonymous):

Yep makes sense :)